OpenStudy (anonymous):

Can someone help me out on finding the amplitude, period, phase shift on sin cos graphs ?

OpenStudy (anonymous):

y=sin(pi-x) i dont now what the period or phase shift is for this equation.

Directrix (directrix):

This may help.

OpenStudy (anonymous):

is the period 2?

jimthompson5910 (jim_thompson5910):

`y=sin(pi-x)` is the same as `y=1*sin(-1x-(-pi))+0` notice how A = 1 B = -1 C = -pi D = 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you just switched the pi-x around?

jimthompson5910 (jim_thompson5910):

yes I wrote it as -x+pi then change that to -x - (-pi)

OpenStudy (anonymous):

period= -2pi?

jimthompson5910 (jim_thompson5910):

more like T = 2pi/|B| T = 2pi/|-1| T = 2pi/1 T = 2pi The period must be positive

OpenStudy (anonymous):

could you teach me how to find the amplitude and period of a given cos or sin graph?

jimthompson5910 (jim_thompson5910):

what is the value of A in this case?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

b=-1

OpenStudy (anonymous):

c=pi

jimthompson5910 (jim_thompson5910):

A = 1 |A| = |1| = 1 is the amplitude

OpenStudy (anonymous):

phase shift=pi

jimthompson5910 (jim_thompson5910):

C = -pi C/B = -pi/(-1) = pi is the phase shift

jimthompson5910 (jim_thompson5910):

D = 0, so y = 0 is the midline

OpenStudy (anonymous):

on my quiz yesterday one of the problems was like...... y=2+3cos(2x-4)......is there a vertical shift in that??

jimthompson5910 (jim_thompson5910):

y=2+3cos(2x-4) is the same as y=3cos(2x-4)+2 D = 2 is the last term tacked on y = 2 is the midline (shifted 2 units up from y = 0)

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

i also had problems on parts where it showed just the graph and no equation, and i had to identify the amplitude, period then the actual equation. can you help me with that. like draw a graph and ask me to identify the period amplitude and equation?

jimthompson5910 (jim_thompson5910):

the amplitude is half the vertical distance from the highest point to the lowest point the period is the horizontal distance from one peak to the next neighboring peak

jimthompson5910 (jim_thompson5910):

wavelength is the same as period http://38.media.tumblr.com/72ab98490685290d67c72d72a859cd93/tumblr_inline_nivaaaEDSf1qi84oh.png

OpenStudy (anonymous):

hmmmm.

OpenStudy (anonymous):

im still having problems

jimthompson5910 (jim_thompson5910):

ok post your image

OpenStudy (anonymous):